# Causality Matters: How Temporal Information Emerges in Video Language Models

### v1

@@ Abstract
Video language models (VideoLMs) have made significant progress in multimodal understanding. However, temporal understanding, which involves identifying event order, duration, and relationships across time, still remains a core challenge. Prior works emphasize positional encodings (PEs) as a key mechanism for encoding temporal structure. Surprisingly, we find that removing or modifying PEs in video inputs yields minimal degradation in the performance of temporal understanding. In contrast, reversing the frame sequence while preserving the original PEs causes a substantial drop. To explain this behavior, we conduct substantial analysis experiments to trace how temporal information is integrated within the model. We uncover a causal information pathway: temporal cues are progressively synthesized through inter-frame attention, aggregated in the final frame, and subsequently integrated into the query tokens. This emergent mechanism shows that temporal reasoning emerges from inter-visual token interactions under the constraints of causal attention, which implicitly encodes temporal structure. Based on these insights, we propose two efficiency-oriented strategies: staged cross-modal attention and a temporal exit mechanism for early token truncation. Experiments on two benchmarks validate the effectiveness of both approaches. To the best of our knowledge, this is the first work to systematically investigate video temporal understanding in VideoLMs, offering insights for future model improvement.
@@ 1 Introduction
Video language models (VideoLMs) (Zhang et al. 2024; Lin et al. 2023), built on the foundation of large language models (LLMs) (OpenAI 2024; Yang et al. 2025), have achieved significant progress in a wide range of video understanding tasks, including captioning, question answering, and temporal reasoning. Among the various challenges posed by these tasks, temporal understanding (Cai et al. 2024; Shangguan et al. 2025), defined as the ability to recognize and interpret the order, duration, and relationships between events over time, stands out as both fundamental and difficult. From reasoning about causal relationships in instructional videos to maintaining narrative coherence in long-form content, temporal understanding plays a vital role in enabling intelligent video-language interactions.
@@
In the pursuit of enhancing the temporal understanding capabilities of VideoLMs (Li et al. 2025; Nguyen et al. 2025), many studies (Liu et al. 2025) have focused on what is likely the model’s most direct method for modeling temporal structures: the positional encodings (PEs). It is commonly assumed that using a more sophisticated positional encoding function leads directly to improved temporal awareness. This belief has spurred a variety of modifications, from extending PEs to higher dimensions (Bai et al. 2025; Wei et al. 2025) to designing variable-rate formats (Ge et al. 2024) to better capture temporal information. However, this intense focus on an explicit signal raises a basic but often overlooked question: To what extent do PEs support temporal understanding in modern VideoLMs?
@@
Based on experiments with the advanced Qwen2.5-VL (Bai et al. 2025) and LLaVA-OneVision (Li et al. 2024), which represent state-of-the-art VideoLMs, we identify several key takeaways about how modern VideoLMs internally process temporal information, as shown in Figure 1:
@@ 2.1 Video Language Model
With the rise of LLMs (OpenAI 2024; Bi et al. 2024), numerous studies (Liu et al. 2024a; Zhang et al. 2024; Yao et al. 2024; Zhang et al. 2025a) have explored their integration into VideoLMs, which increasingly outperform traditional vision-only methods. One major direction focuses on enhancing the video encoder (Zhao et al. 2024b; Choudhury et al. 2024; Chung et al. 2025). Another line of work (Chen et al. 2024; Wei et al. 2025; Ge et al. 2024; Bai et al. 2025) adapts LLMs to better accommodate multimodal inputs. In addition, other studies aim to improve the efficiency of processing long video inputs (Fu et al. 2024; He et al. 2024; Shi, Long, and Wang 2025). Collectively, these advancements have driven progress in video-language understanding.
@@ 2.2 Temporal Understanding
VideoLMs face unique challenges in video temporal understanding, which requires detecting similarities and differences across frames. Benchmarks (Cai et al. 2024; Shangguan et al. 2025; Plizzari et al. 2025; Liu et al. 2024b) have been proposed to evaluate this capability. Besides, various approaches (Nie et al. 2024; Hu et al. 2024; Zhao et al. 2025; Fateh et al. 2024) have been introduced to improve this capability. Among them, T3 (Li et al. 2025) transfers temporal skills from synthetic text tasks to VideoLMs. Despite these advances, the internal mechanisms by which VideoLMs capture temporal information remain underexplored. Understanding how temporal information is extracted is essential for building more robust and generalizable models.
@@ 2.3 Model Mechanistic Interpretability
Many studies analyze how LLMs make predictions (Zhao et al. 2024a; Goldshmidt and Horovicz 2024; Biran et al. 2024; Wang et al. 2023). This line of research has also been extended to multimodal large language models (Yu and Ananiadou 2024; Golovanevsky et al. 2025; Zhang et al. 2025b). For instance, Basu et al. (2024) uses causal tracing to reveal how early-layer modules capture and transmit visual information. However, most existing work focuses on single-image tasks, and the interpretability of VideoLMs, particularly for temporal reasoning, remains underexplored. This gap motivates our work, which seeks to explain how temporal information is captured within VideoLMs.
@@ 3.1 Dataset
We conduct analysis on TempCompass (Liu et al. 2024b), which encompasses a diverse range of temporal phenomena, including action, speed, direction, attribute change, and event order. It provides a fine-grained assessment of a model’s ability to reason over temporal information. The benchmark includes four task formats, summarized in Table 1. Since our evaluation focuses on the model’s first predicted token, we convert the caption generation task into a multiple-choice format, where the model is required to select the most temporally accurate option.
To validate our proposed efficiency-oriented strategies, we run downstream application experiments on the multiple-choice subset of the NExT-QA dataset (Xiao et al. 2021), which emphasizes explanatory reasoning over causal and temporal relationships. The test set contains 8,564 questions. To further assess generalizability, we additionally report results on the open-ended ActivityNet-QA dataset (Caba Heilbron et al. 2015), included in the appendix.
@@ 3.2 Model
We evaluate two representative VideoLMs: Qwen2.5-VL and LLaVA-OneVision. Qwen2.5-VL (Bai et al. 2025) uses multi-dimensional positional encodings tailored for video, making it well-suited to study PE impact on spatiotemporal modeling. In contrast, LLaVA-OneVision (Li et al. 2024) is another state-of-the-art multimodal model that does not adapt PEs for video, allowing us to derive insights that are agnostic to special PE designs. Unless specialized otherwise, we report results using Qwen2.5-VL-7B in the main text. Results for Qwen2.5-VL-3B and LLaVA-OneVision are provided in the appendix.
@@ 5 Temporal Information Emerges from Attention-Based Causal Interactions
To investigate this hypothesis, we examine how causal attention mechanisms facilitate the step-by-step construction and propagation of temporal information throughout the model. Drawing inspiration from Zhang et al. (2025b), we adopt a backward tracing approach, analyzing the model from its output back to earlier interactions. We identify where temporal information originates, how it flows between video and text modalities, and validate its causal nature through targeted analysis.
We adapt the attention knockout method (Geva et al. 2023) to the VideoLM setting in order to analyze this process. Specifically, we modify the attention computation in Equation 2 by incorporating an additional mask $\bm{M}$ that prevents target tokens in the set $\mathcal{T}$ from attending to source tokens in the set $\mathcal{S}$:

### v2

@@ Abstract
Video language models (VideoLMs) have made significant progress in multimodal understanding. However, temporal understanding, which involves identifying event order, duration, and relationships across time, still remains a core challenge. Prior works emphasize positional encodings (PEs) as a key mechanism for encoding temporal structure. Surprisingly, we find that removing or modifying PEs in video inputs yields minimal degradation in the performance of temporal understanding. In contrast, reversing the frame sequence while preserving the original PEs causes a substantial drop. To explain this behavior, we conduct substantial analysis experiments to trace how temporal information is integrated within the model. We uncover a causal information pathway: temporal cues are progressively synthesized through inter-frame attention, aggregated in the final frame, and subsequently integrated into the query tokens. This emergent mechanism shows that temporal reasoning emerges from inter-visual token interactions under the constraints of causal attention, which implicitly encodes temporal structure. Based on these insights, we propose two efficiency-oriented strategies: staged cross-modal attention and a temporal exit mechanism for early token truncation. Experiments on two benchmarks validate the effectiveness of both approaches. To the best of our knowledge, this is the first systematic study of video temporal understanding in VideoLMs, offering insights for future model improvement. Our code is available at https://github.com/ANDgate99/Causality-Matters.
@@ 1 Introduction
Video language models (VideoLMs) (zhang2024longva; lin2023video), built on the foundation of large language models (LLMs) (openai2024gpt4o; yang2025qwen3), have achieved significant progress in a wide range of video understanding tasks, including captioning, question answering, and temporal reasoning. Among the various challenges posed by these tasks, temporal understanding (cai2024temporalbench; shangguantomato), defined as the ability to recognize and interpret the order, duration, and relationships between events over time, stands out as both fundamental and difficult. From reasoning about causal relationships in instructional videos to maintaining narrative coherence in long-form content, temporal understanding plays a vital role in enabling intelligent video-language interactions.
@@
In the pursuit of enhancing the temporal understanding capabilities of VideoLMs (litemporal; nguyen2025moose), many studies (liu2025vrope) have focused on what is likely the model’s most direct method for modeling temporal structures: the positional encodings (PEs). It is commonly assumed that using a more sophisticated positional encoding function leads directly to improved temporal awareness. This belief has spurred a variety of modifications, from extending PEs to higher dimensions (Qwen2.5-VL; weivideorope) to designing variable-rate formats (ge2024v2pe) to better capture temporal information. However, this intense focus on an explicit signal raises a basic but often overlooked question: To what extent do PEs support temporal understanding in modern VideoLMs?
@@
Based on experiments with the advanced Qwen2.5-VL (Qwen2.5-VL) and LLaVA-OneVision (li2024llava), which represent state-of-the-art VideoLMs, we identify several key takeaways about how modern VideoLMs internally process temporal information, as shown in Figure 1:
@@ 2.1 Video Language Model
With the rise of LLMs (openai2024gpt4o; bi2024deepseek), numerous studies (kangaroogroup; zhang2024longva; yao2024minicpm; damonlpsg2025videollama3) have explored their integration into VideoLMs, which increasingly outperform traditional vision-only methods. One major direction focuses on enhancing the video encoder (zhao2024videoprism; choudhury2024don; chungunifying). Another line of work (chen2024internvl; weivideorope; ge2024v2pe; Qwen2.5-VL) adapts LLMs to better accommodate multimodal inputs. In addition, other studies aim to improve the efficiency of processing long video inputs (fu2024framefusion; he2024ma; shi2025static). Collectively, these advancements have driven progress in video-language understanding.
@@ 2.2 Temporal Understanding
VideoLMs face unique challenges in video temporal understanding, which requires detecting similarities and differences across frames. Benchmarks (cai2024temporalbench; shangguantomato; plizzari2025omnia; liu2024tempcompass) have been proposed to evaluate this capability. Besides, various approaches (nie2024slowfocus; hu2024enhancing; zhao2025videoexpert; fateh2024video) have been introduced to improve this capability. Among them, T3 (litemporal) transfers temporal skills from synthetic text tasks to VideoLMs. Despite these advances, the internal mechanisms by which VideoLMs capture temporal information remain underexplored. Understanding how temporal information is extracted is essential for building more robust and generalizable models.
@@ 2.3 Model Mechanistic Interpretability
Many studies analyze how LLMs make predictions (zhao2024towards; goldshmidt2024tokenshap; biran2024hopping; wang2023label). This line of research has also been extended to multimodal large language models (yu2024understanding; golovanevsky2025vlms; zhang2025cross). For instance, basu2024understanding uses causal tracing to reveal how early-layer modules capture and transmit visual information. However, most existing work focuses on single-image tasks, and the interpretability of VideoLMs, particularly for temporal reasoning, remains underexplored. This gap motivates our work, which seeks to explain how temporal information is captured within VideoLMs.
@@ 3.1 Dataset
We conduct analysis on TempCompass (liu2024tempcompass), which encompasses a diverse range of temporal phenomena, including action, speed, direction, attribute change, and event order. It provides a fine-grained assessment of a model’s ability to reason over temporal information. The benchmark includes four task formats, summarized in Table 1. Since our evaluation focuses on the model’s first predicted token, we convert the caption generation task into a multiple-choice format, where the model is required to select the most temporally accurate option.
To validate our proposed efficiency-oriented strategies, we run downstream application experiments on the multiple-choice subset of the NExT-QA dataset (xiao2021next), which emphasizes explanatory reasoning over causal and temporal relationships. The test set contains 8,564 questions. To further assess generalizability, we additionally report results on the open-ended ActivityNet-QA dataset (caba2015activitynet), included in the appendix.
@@ 3.2 Model
We evaluate two representative VideoLMs: Qwen2.5-VL and LLaVA-OneVision. Qwen2.5-VL (Qwen2.5-VL) uses multi-dimensional positional encodings tailored for video, making it well-suited to study PE impact on spatiotemporal modeling. In contrast, LLaVA-OneVision (li2024llava) is another state-of-the-art multimodal model that does not adapt PEs for video, allowing us to derive insights that are agnostic to special PE designs. Unless specialized otherwise, we report results using Qwen2.5-VL-7B in the main text. Results for Qwen2.5-VL-3B and LLaVA-OneVision are provided in the appendix.
@@ 5 Temporal Information Emerges from Attention-Based Causal Interactions
To investigate this hypothesis, we examine how causal attention mechanisms facilitate the step-by-step construction and propagation of temporal information throughout the model. Drawing inspiration from zhang2025cross, we adopt a backward tracing approach, analyzing the model from its output back to earlier interactions. We identify where temporal information originates, how it flows between video and text modalities, and validate its causal nature through targeted analysis.
We adapt the attention knockout method (geva2023dissecting) to the VideoLM setting in order to analyze this process. Specifically, we modify the attention computation in Equation 2 by incorporating an additional mask $\bm{M}$ that prevents target tokens in the set $\mathcal{T}$ from attending to source tokens in the set $\mathcal{S}$:
@@ 7 Conclusion
Acknowledgment
This research is supported by the NTU Start-Up Grant (#023284-00001), Singapore, and the MOE AcRF Tier 1 Seed Funding Grant (#025041-00001, RS37/24).
Appendix
The appendix presents extended results from our analysis using Qwen2.5-VL-3B and LLaVA-OneVision-7B to demonstrate the generalizability of our findings. Additionally, we include results on open-ended questions from the ActivityNet-QA dataset to assess the downstream effectiveness of strategies based on our earlier findings.
Appendix A Additional Analysis Results
A.1 Qwen2.5-VL-3B
The Qwen2.5-VL-3B model and the reported Qwen2.5-VL-7B in the main text belong to the same model family and share fundamental architectural similarities, particularly in their positional encoding designs and multimodal input processing mechanisms. However, they exhibit distinct structural configurations: while the 7B variant employs a wider but shallower architecture with 28 layers, 28 attention heads, and a hidden size of 3,584, the 3B model adopts a narrower yet deeper structure comprising 36 layers, 16 attention heads, and a 2,048-dimensional hidden space. This structural variation allows us to assess how key findings generalize across different depth–width configurations in video language models, with the present study focusing specifically on the 3B variant to evaluate the robustness of the observed phenomena.
Figure 10: Layer-wise PE ablation effects on Qwen2.5-VL-3B. As with the 7B variant, ablating positional encodings significantly affects only the first layer, with negligible impact on subsequent layers.
Analyzing the Impact of Positional Encodings.
We conduct the same positional encoding (PE) ablation analysis on the Qwen2.5-VL-3B model as described in the main text. As shown in Figure 10, the impact of removing PEs is largely confined to the first layer, with consistent trends observed across multiple task types. This pattern closely mirrors the behavior seen in the Qwen2.5-VL-7B model. Interestingly, the 3B variant exhibits more uniform performance across task types, possibly due to its deeper architecture facilitating more consistent information propagation and representation learning across layers. These findings suggest that, under different architectural regimes (e.g., deep-narrow vs. shallow-wide), PEs in most layers contribute minimally to temporal understanding.
Figure 11: Impact of modality-specific position ID shuffling at the first layer in Qwen2.5-VL-3B. Disrupting positional IDs for textual inputs significantly degrades temporal understanding performance, while doing so for video inputs has minimal effect.
As shown in Figure 11, in the modality-specific position ID shuffling experiment, the Qwen2.5-VL-3B model exhibits behavior highly consistent with that of the 7B variant: the observed impact on temporal understanding primarily stems from shuffling PEs of textual inputs, while disrupting video PEs has negligible effects. However, overall, the effect of textual PEs in the 3B model appears weaker than in the 7B model. This difference may be attributed to the deeper architecture of the 3B variant, which could allow greater flexibility in compensating for disrupted positional signals through deeper contextual integration.
Figure 12: Effect of reversing position IDs versus frame order in Qwen2.5-VL-3B. Reversing the frame order causes a pronounced drop in ground-truth probability, while reversing PEs produces only a minimal impact, further demonstrating their limited effectiveness.
In the experiment comparing the effects of ablating PEs and altering frame order, we observe that reversing the position IDs of video frames leads to only a minor drop in ground-truth probability within 1.4% across all task types in Figure 12. In contrast, reversing the frame order while preserving the original position IDs results in substantial degradation, with performance drops of up to 8.93%. These findings underscore the critical role of frame ordering in temporal understanding, while further highlighting the limited contribution of PEs across model structures.
Figure 13: Effect of blocking attention from the final token to different input modalities in Qwen2.5-VL-3B. Restricting attention to query tokens substantially degrades temporal understanding performance, while blocking video tokens has a negligible effect.
Exploring Causal Temporal Generation Pathway.
As shown in Figure 13, blocking attention from the final prediction token to video tokens in Qwen2.5-VL-3B continues to have minimal impact. In contrast, restricting access to query tokens results in significant performance drops, particularly in deeper layers. Interestingly, unlike the 7B model, the 3B variant also shows sensitivity to query blocking in the first few layers, suggesting that it relies on query tokens earlier in the processing pipeline. This early drop reflects a structural adaptation in the deeper, narrower architecture of the 3B model, which enables earlier layers to more effectively extract and utilize query information. Nonetheless, across both models, query tokens remain the dominant source of temporal cues for final-stage prediction.
Figure 14: Effect of blocking inter-frame and frame-to-query attention in Qwen2.5-VL-3B. Inter-frame blocking causes early ground-truth probability drops, marking temporal encoding onset; frame-to-query blocking causes later drops, showing information flow to the query.
Figure 14 shows that Qwen2.5-VL-3B also exhibits a two-stage temporal processing pattern: early drops from inter-frame attention blocking indicate where temporal features are composed across frames, while later drops from frame-to-query blocking reveal when this information flows to the query. Compared to the 7B model, the inter-frame effects in 3B are notably milder, suggesting that the deeper architecture of the 3B model may be less effective at capturing temporal information through inter-frame interactions, an observation consistent with intuition.
Figure 15: Effect of restricting query attention to single frames in Qwen2.5-VL-3B. Attending only to early frames leads to greater performance drops, suggesting limited temporal information accumulated at those points; minimal impact when attending to the last frame indicates that temporal information has largely converged there.
In Qwen2.5-VL-3B, the final frame also appears to aggregate most of the temporal information, as Figure 15 shows that restricting query attention to the last frame in middle layers results in minimal performance drop, typically under 2%. However, the differences between attending to earlier frames (e.g., Frame 1 vs. Frame 2) are less pronounced compared to the 7B model, suggesting that Qwen2.5-VL-3B has a weaker capacity for inter-frame temporal aggregation, likely due to its smaller model size.
Figure 16: Effect of reversing video order in the “Yes or No” task in Qwen2.5-VL-3B. Attribution patterns invert, suggesting that temporal understanding depends on causal processing rather than simple visual accumulation.
As shown in Figure 16, for the “Yes or No” task, reversing the frame order completely flips the attribution: forcing the query to attend to the last frame now yields the worst performance, even when the position IDs still follow the original frame order. This indicates that temporal information in the 3B model is also aggregated in a causal, order-sensitive manner, rather than through simple accumulation.
Figure 17: Effect of restricting attention to analyze inter-frame spatiotemporal contributions in Qwen2.5-VL-3B. Sparse, long-range attention outperforms dense, local attention. Spatially, early layers favor global context, while later layers attend to finer details.
Qwen2.5-VL-3B and the 7B model share a similar process for constructing inter-frame temporal information. As shown in Figure 17, restricting each frame to attend only to corresponding regions across all previous frames yields better performance than attending to the entire previous frame, despite involving fewer total token interactions. From a temporal perspective, the length of context appears more important than its density. Spatially, the difference between attending to the full previous frame and its corresponding region is evident only in the early and middle layers, with little distinction in deeper layers. This suggests that the model initially relies on global spatial context and gradually shifts its focus to finer local details as depth increases.
A.2 LLaVA-OneVision-7B
LLaVA-OneVision-7B differs from the Qwen2.5-VL model series in several architectural aspects. For instance, it does not merge video frame features. However, a more critical distinction is its lack of any dedicated positional encoding design for video inputs. Instead, it directly inherits the rotary positional encoding from its base language model, applying simple 1D position IDs sequentially across video tokens, without explicitly modeling frame boundaries or spatial structure. The use of this standard 1D encoding, which has proven effective in large language models, helps eliminate the possibility that the ineffectiveness of positional encodings is due to overly complex or poorly tuned design choices. Therefore, this minimal design makes LLaVA-OneVision-7B a strong candidate for evaluating the generalizability and validity of our findings on positional encodings.
Figure 18: Layer-specific effects of removing positional encodings in LLaVA-OneVision-7B. The removal of positional encodings only has a strong effect on the first layer but minimal impact on later layers.
As shown in Figure 18, the layer-wise PE ablation results on LLaVA-OneVision-7B follow the same trend observed in Qwen2.5-VL models. Removing PEs in the first layer leads to a substantial drop in the prediction probability of the ground-truth, exceeding 60% in nearly all task types. In contrast, removing PEs from subsequent layers results in negligible performance change, with the ground-truth probability curve remaining nearly flat throughout. This consistent pattern, despite architectural and design differences across models, suggests that positional encodings primarily contribute to temporal understanding only in the first layer, while later layers rely more on learned contextual features rather than explicit position signals.
Figure 19: Effect of modality-specific position ID shuffling at the first layer in LLaVA-OneVision-7B. Altering positional IDs for textual inputs leads to a significant drop in temporal understanding performance, whereas a similar disruption for video inputs has little impact.
As shown in Figure 19, LLaVA-OneVision-7B exhibits a similar trend in the modality-specific position ID shuffling experiment. Shuffling the position IDs of textual tokens at the first layer results in a substantial drop in ground-truth prediction probability across all tasks, up to 58.88% for “Caption Matching” task. In contrast, shuffling the position IDs of video tokens has minimal impact, with performance changes remaining below 0.1%. These results indicate that even simple and effective 1D positional encodings, when applied to video inputs, still fail to contribute meaningfully to temporal information capture.
Figure 20: Effect of reversing position IDs versus frame order in LLaVA-OneVision-7B. Reversing frame order still causes a clear drop, while reversing position IDs has minimal effect. Compared to Qwen2.5-VL, the gap is smaller, suggesting that simpler designs may be more easily utilized.
We further evaluate the role of PEs in temporal modeling by comparing the effects of reversing frame order versus reversing position IDs in LLaVA-OneVision-7B. As shown in Figure 20, reversing the frame order leads to a noticeable drop in ground-truth prediction probability, indicating that the model is still sensitive to the frame order. In contrast, reversing the position IDs results in only minor performance degradation, with changes mostly within 4%. This trend aligns with findings in Qwen2.5-VL models, reaffirming the limited impact of PEs. However, the performance gap between the two conditions is smaller in LLaVA-OneVision-7B, suggesting that its simpler PE design may be more easily exploited by the model.
Appendix B Additional Downstream Task Verification
We evaluate three proposed efficiency strategies on the ActivityNet-QA dataset using the Qwen2.5-VL-7B model. The strategies are as follows: (1) between layers 10–20, restrict the query’s attention to only the final frame in order to reduce computation; (2) disable inter-frame attention across layers 20–28 to reduce computation; and (3) completely discard all frame tokens after layer 20 to reduce memory usage. ActivityNet-QA is an open-ended question-answering task, so we allow the model to generate responses of up to 100 tokens per question. The quality of generated answers is assessed using GPT-4o-mini as the evaluator.
Metric Baseline (1) (2) (3)
Accuracy 46.9 40.4 46.9 42.7
Score 2.71 2.39 2.71 2.54
Table 3: Accuracy and score on ActivityNet-QA across different efficient inference strategies based on our findings.
As shown in Table 3, Strategy (2), which disables inter-frame attention between layers 20–28, achieves the same accuracy (46.9%) and score (2.71) as the full baseline model, demonstrating that later-stage inter-frame interactions can be safely pruned without sacrificing any performance. Strategy (3), which discards all frame tokens after layer 20, shows a moderate drop in accuracy (42.7%), suggesting that much of the temporal understanding is already completed in earlier layers. It is worth noting that ActivityNet-QA is a more challenging dataset, with longer video durations and a stronger reliance on temporal information. Strategy (1), which limits query attention to only the final frame between layers 10–20, results in a more noticeable performance degradation. Despite being relatively aggressive, the strategy maintains over 40% accuracy, suggesting potential for further optimization to reduce performance loss. Overall, these results validate the practicality and insightfulness of our findings and highlight their potential for enabling more efficient model designs.